# Towards Massive Interaction with Generalist Robotics: A Systematic Review of XR-enabled Remote Human-Robot Interaction Systems

### v2

@@ Abstract.
This survey provides an exhaustive review of the applications of extended reality (XR) technologies in the field of remote human-computer interaction (HRI). We developed a systematic search strategy based on the PRISMA methodology. From the initial 2,561 articles selected, 100 research papers that met our inclusion criteria were included. We categorized and summarized the domain in detail, delving into XR technologies, including augmented reality (AR), virtual reality (VR), and mixed reality (MR), and their applications in facilitating intuitive and effective remote control and interaction with robotic systems. The survey highlights existing articles on the application of XR technologies, user experience enhancement, and various interaction designs for XR in remote HRI, providing insights into current trends and future directions. We also identified potential gaps and opportunities for future research to improve remote HRI systems through XR technology to guide and inform future XR and robotics research.
@@ 2.2. Inclusion and Exclusion Criteria
Criterion
Description
$I_{1}$
The research proposes the design, development or system for remote control of robots based on augmented, virtual or mixed reality.
$I_{2}$
The research used augmented, virtual or mixed reality as a method of remote control of the robot.
$E_{1}$
The system proposed in the study does not support remote control.
$E_{2}$
The study does not involve real robots.
$E_{3}$
The target of research control is not a robot.
$E_{4}$
The study did not use augmented, virtual or mixed reality technologies.
$E_{5}$
This is not a technical article.
@@ 2.3. Data Extraction and Analysis
ID
Data Extraction
Type
DE1
Study ID
Open text
DE2
Title
Open text
DE3
Keywords
Open text
DE4
Used XR technologies
VR HMD, AR HMD, MR, Mobile AR, CAVE, Other
DE5
Interaction modalities
Gesture, Controller, Joystick, Gaze, Head, Haptic devices, Motion capture, Walk, 2D screen, Voice, Glove, Other
DE6
Virtual interfaces
Direct, Digital twin, Virtual control room, Digital twin+3D reconstruction, Direct+3D reconstruction, 3D reconstruction, Context-aware AR, Multiple, Other
DE7
User perspective
Coupling with robots, Decoupling from robots, Dynamic perspective, Bird’s-eye view, Other
DE8
Type of robots
Mobile robot, Drones/UAV, Humanoid Robot, Robotic arm, Mobile robot+robotic arm, Double-armed robot, Medical Robotics, Other
DE9
Specific tasks involved
Navigation, Grabbing/Picking/Placement, Surgery, Game/Entertainment, Industrial/Manufacturing, Search, Environment scan, No, Multiple, Other
DE10
How was it measured or evaluated?
Time/accuracy of the task, Interviews, Questionnaire, AR/VR performance, Comparison, N/A, Other
DE11
Was there a discussion about delays?
No, Yes(Times), Other
DE12
Support multiplayer collaboration?
No, Multi-user - one robot, One user - multi-robot, Multi - Multi, Other
DE13
Where are the enhancements located?
User, Robot (real), Robot (virtual), Object (virtual), Real environment(RE), Virtual environment(VE)
DE14
Enhancement types
Haptic, Voice, Graphic, Text, 3D Object, Highlight, Ray, Avatar
@@ Appendix A Data Extraction List for Included Publication
VR HMD
(Chen et al., 2022)(Xu et al., 2022b)(Ponomareva et al., 2021)(Wei et al., 2021)(Vu et al., 2022)(Lipton et al., 2017)(Su et al., 2021)(Walker et al., 2021)(Wibowo et al., 2021)(Zein et al., 2021)(Vagvolgyi et al., 2018)(Kulikovskiy et al., 2021)(Zhou et al., 2020)(Chen et al., 2017)(Park et al., 2018)(Stotko et al., 2019)(Peppoloni et al., 2015b)(Wang et al., 2019a)(Chen et al., 2020)(Galambos et al., 2015)(Kuo et al., 2021)(Yun and Jun, 2022)(Xie et al., 2022)(Schwarz et al., 2021)(Cichon and Roßmann, 2018)(Vempati et al., 2019)(Fani et al., 2018b)(Zinchenko and Song, 2021)(Codd-Downey et al., 2014)(Bai et al., 2018)(Peppoloni et al., 2015a)
@@
AR HMD
(Mourtzis et al., 2019)(Hedayati et al., 2018)(Mourtzis et al., 2017)(Xie et al., 2022)(Hernández et al., 2020)(Gai et al., 2020)(Du et al., 2022b)(Walker et al., 2019)(Xue Er Shamaine et al., 2020)(Du et al., 2019)(Choi et al., 2022)(Gong et al., 2017)(Xue et al., 2020)(Aschenbrenner et al., 2018)(Szczurek et al., 2023)(Steinke et al., 2023)(Meng et al., 2023)(Rivera-Pinto et al., 2023)(Black et al., 2023)(Du et al., 2022a)
MR
(Cruz Ulloa et al., 2022)(Ai et al., 2016)(Hönig et al., 2015)(Ihara et al., 2023)(Zaman et al., 2023)
CAVE
(Galambos et al., 2015)(Gammieri et al., 2017)(Betancourt et al., 2022)
Mobie AR
(Chacko et al., 2020)
Other
(Stedman et al., 2022)(Rastogi and Kumar Srivastava, 2019)(Yew et al., 2017)(Mu et al., 2021)
@@
Controller
(Chen et al., 2022)(Wei et al., 2021)(Lipton et al., 2017)(Walker et al., 2021)(Kulikovskiy et al., 2021)(Zhou et al., 2020)(Wang et al., 2019a)(Chen et al., 2020)(Gammieri et al., 2017)(Kuo et al., 2021)(Vempati et al., 2019)(Codd-Downey et al., 2014)(Takacs et al., 2015)(Omarali et al., 2020)(Xu et al., 2022a)(Kalinov et al., 2021)(Hernández et al., 2020)(Le et al., 2020)(Wang et al., 2017)(Nagy et al., 2019)(Nandhini et al., 2023)(Abdulsalam and Aouf, 2023)(Kanazawa et al., 2023)(Fuchino et al., 2023)(Nenna et al., 2023)(Sanfilippo et al., 2022)(Yun and Jun, 2022)(Zaman et al., 2023)(Fan et al., 2023)(Liu et al., 2023)
Gesture
(Cruz Ulloa et al., 2022)(Mourtzis et al., 2017)(Xie et al., 2022)(Peppoloni et al., 2015b)(Jang et al., 2021a)(Jang et al., 2021b)(Zhao et al., 2017)(Xu et al., 2018)(Cousins et al., 2017)(Xue Er Shamaine et al., 2020)(Bian et al., 2018)(Choi et al., 2022)(Xue et al., 2020)(Ihara et al., 2023)(Szczurek et al., 2023)(Meng et al., 2023)(Black et al., 2023)(Du et al., 2022b)(Sakashita et al., 2023)(Zaman et al., 2023)(Rivera-Pinto et al., 2023)
Joystick
(Vu et al., 2022)(Wibowo et al., 2021)(Zein et al., 2021)(Hedayati et al., 2018)(Stotko et al., 2019)(Ai et al., 2016)(Betancourt et al., 2022)(Walker et al., 2019)(Xia et al., 2023)
Haptic devices
(Ponomareva et al., 2021)(Chen et al., 2017)(Su et al., 2021)(Rastogi and Kumar Srivastava, 2019)(Schwarz et al., 2021)(Trejo and Hu, 2018)(Kim et al., 2023)(Yun and Jun, 2022)(Fan et al., 2023)
Motion capture
(Xu et al., 2022b)(Peppoloni et al., 2015b)(Fani et al., 2018b)(Bai et al., 2018)(Brizzi et al., 2018)
Walk
(Phan et al., 2018)(Gai et al., 2020)(Grzeskowiak et al., 2020)(Du et al., 2019)(Sakashita et al., 2023)
2D screen
(Mu et al., 2021)(Chacko et al., 2020)(Gong et al., 2017)
Voice
(Hormaza et al., 2019)(Aschenbrenner et al., 2018)(Rivera-Pinto et al., 2023)
Glove
(Park et al., 2018)(Sun et al., 2021)(Trinitatova and Tsetserukou, 2023)
Head
(Theofilis et al., 2016)(Sakashita et al., 2023)
Gaze
(Moniri et al., 2016)
Other
(Stedman et al., 2022)(Mourtzis et al., 2019)(Vagvolgyi et al., 2018)(Galambos et al., 2015)(Cichon and Roßmann, 2018)(Zinchenko and Song, 2021)(Wang et al., 2019c)(Hönig et al., 2015)(Wang et al., 2021b)(Du et al., 2019)(McHenry et al., 2021)(Liu et al., 2023)
@@
Digital twin
(Cruz Ulloa et al., 2022)(Su et al., 2021)(Vagvolgyi et al., 2018)(Wang et al., 2019a)(Yew et al., 2017)(Gammieri et al., 2017)(Yun and Jun, 2022)(Xie et al., 2022)(Zinchenko and Song, 2021)(Peppoloni et al., 2015b)(Takacs et al., 2015)(Jang et al., 2021b)(Trejo and Hu, 2018)(Betancourt et al., 2022)(Le et al., 2020)(Wang et al., 2021b)(Du et al., 2022b)(Walker et al., 2019)(Brizzi et al., 2018)(Xue Er Shamaine et al., 2020)(Grzeskowiak et al., 2020)(Wang et al., 2019b)(Nagy et al., 2019)(Choi et al., 2022)(McHenry et al., 2021)(Xue et al., 2020)(Nandhini et al., 2023)(Fan et al., 2023)(Abdulsalam and Aouf, 2023)(Trinitatova and Tsetserukou, 2023)(Meng et al., 2023)
@@
Direct
(Xu et al., 2022b)(Wibowo et al., 2021)(Hedayati et al., 2018)(Hedayati et al., 2018)(Chen et al., 2017)(Park et al., 2018)(Stotko et al., 2019)(Peppoloni et al., 2015b)(Chen et al., 2020)(Ai et al., 2016)(Schwarz et al., 2021)(Fani et al., 2018b)(Bai et al., 2018)(Wang et al., 2019c)(Mu et al., 2021)(Theofilis et al., 2016)(Zhao et al., 2017)(Gai et al., 2020)(Cousins et al., 2017)(Chacko et al., 2020)(Brizzi et al., 2018)(Bian et al., 2018)(Zhang, 2018)(Gong et al., 2017)(Ihara et al., 2023)(Zaman et al., 2023)(Xia et al., 2023)(Black et al., 2023)(Sanfilippo et al., 2022)
Digital twin+3D reconstruction
(Ponomareva et al., 2021)(Stedman et al., 2022)(Rastogi and Kumar Srivastava, 2019)(Galambos et al., 2015)(Kuo et al., 2021)(Cichon and Roßmann, 2018)(Codd-Downey et al., 2014)(Phan et al., 2018)(Omarali et al., 2020)(Jang et al., 2021a)(Hönig et al., 2015)(Wang et al., 2017)(Szczurek et al., 2023)(Du et al., 2022b)(Aschenbrenner et al., 2018)
Multiple
(Vu et al., 2022)(Zhou et al., 2020)(Xu et al., 2022a)(Du et al., 2019)(Kanazawa et al., 2023)(Liu et al., 2023)
Direct+3D reconstruction
(Chen et al., 2022)(Zein et al., 2021)(Vempati et al., 2019)(Sakashita et al., 2023)(Kim et al., 2023)
3D reconstruction
(Wei et al., 2021)(Walker et al., 2021)(Moniri et al., 2016)
Virtual control room
(Lipton et al., 2017)(Kulikovskiy et al., 2021)(Steinke et al., 2023)
Other
(Mourtzis et al., 2019)(Mourtzis et al., 2017)(Hernández et al., 2020)(Sun et al., 2021)(Xu et al., 2018)(Hormaza et al., 2019)(Fuchino et al., 2023)
@@
Decoupling with robots
(Ponomareva et al., 2021)(Vu et al., 2022)(Stedman et al., 2022)(Cruz Ulloa et al., 2022)(Mourtzis et al., 2019)(Su et al., 2021)(Hedayati et al., 2018)(Vagvolgyi et al., 2018)(Wang et al., 2019a)(Rastogi and Kumar Srivastava, 2019)(Yew et al., 2017)(Galambos et al., 2015)(Gammieri et al., 2017)(Kuo et al., 2021)(Yun and Jun, 2022)(Xie et al., 2022)(Cichon and Roßmann, 2018)(Zinchenko and Song, 2021)(Codd-Downey et al., 2014)(Peppoloni et al., 2015a)(Omarali et al., 2020)(Jang et al., 2021b)(Trejo and Hu, 2018)(Moniri et al., 2016)(Xu et al., 2018)(Le et al., 2020)(Hönig et al., 2015)(Wang et al., 2021b)(Du et al., 2022b)(Hormaza et al., 2019)(Walker et al., 2019)
@@
Coupling with robots
(Xu et al., 2022b)(Lipton et al., 2017)(Wibowo et al., 2021)(Zein et al., 2021)(Kulikovskiy et al., 2021)(Chen et al., 2017)(Park et al., 2018)(Stotko et al., 2019)(Peppoloni et al., 2015b)(Mourtzis et al., 2017)(Chen et al., 2020)(Ai et al., 2016)(Schwarz et al., 2021)(Vempati et al., 2019)(Fani et al., 2018b)(Bai et al., 2018)(Wang et al., 2019c)(Mu et al., 2021)(Kalinov et al., 2021)(Theofilis et al., 2016)(Zhao et al., 2017)(Gai et al., 2020)(Sun et al., 2021)(Cousins et al., 2017)(Chacko et al., 2020)(Brizzi et al., 2018)(Bian et al., 2018)(Zhang, 2018)(Gong et al., 2017)(Sakashita et al., 2023)(Kim et al., 2023)(Steinke et al., 2023)
@@
Dynamic perspective
(Zhou et al., 2020)(Xu et al., 2022a)(Hernández et al., 2020)(Betancourt et al., 2022)(Liu et al., 2023)
God perspective
(Walker et al., 2021)(Jang et al., 2021a)
Other
(Chen et al., 2022)(Vu et al., 2022)(Takacs et al., 2015)(Phan et al., 2018)(Grzeskowiak et al., 2020)(Aschenbrenner et al., 2018)(Fuchino et al., 2023)(Black et al., 2023)
@@
Robotic arm
(Xu et al., 2022b)(Ponomareva et al., 2021)(Wei et al., 2021)(Vu et al., 2022)(Su et al., 2021)(Vagvolgyi et al., 2018)(Chen et al., 2017)(Wang et al., 2019a)(Rastogi and Kumar Srivastava, 2019)(Galambos et al., 2015)(Gammieri et al., 2017)(Kuo et al., 2021)(Yun and Jun, 2022)(Fani et al., 2018b)(Bai et al., 2018)(Peppoloni et al., 2015a)(Omarali et al., 2020)(Xu et al., 2022a)(Jang et al., 2021b)(Xu et al., 2018)(Le et al., 2020)(Cousins et al., 2017)(Wang et al., 2021b)(Hormaza et al., 2019)(Xue Er Shamaine et al., 2020)(Wang et al., 2019b)(Du et al., 2019)(Nagy et al., 2019)(Choi et al., 2022)(McHenry et al., 2021)
@@
Mobile robot
(Stedman et al., 2022)(Cruz Ulloa et al., 2022)(Walker et al., 2021)(Wibowo et al., 2021)(Stotko et al., 2019)(Cichon and Roßmann, 2018)(Codd-Downey et al., 2014)(Jang et al., 2021a)(Hernández et al., 2020)(Zhao et al., 2017)(Gai et al., 2020)(Du et al., 2022b)(Chacko et al., 2020)(Wang et al., 2017)(Zhang, 2018)(Ihara et al., 2023)(Sakashita et al., 2023)(Szczurek et al., 2023)(Kanazawa et al., 2023)(Xia et al., 2023)
Drones/UAV
(Zein et al., 2021)(Hedayati et al., 2018)(Ai et al., 2016)(Vempati et al., 2019)(Takacs et al., 2015)(Wang et al., 2019c)(Kalinov et al., 2021)(Betancourt et al., 2022)(Hönig et al., 2015)(Walker et al., 2019)
Humanoid robot
(Chen et al., 2022)(Kulikovskiy et al., 2021)(Schwarz et al., 2021)(Takacs et al., 2015)(Theofilis et al., 2016)(Grzeskowiak et al., 2020)(Gong et al., 2017)(Sanfilippo et al., 2022)
Double-armed robot
(Lipton et al., 2017)(Zhou et al., 2020)(Park et al., 2018)(Peppoloni et al., 2015b)(Moniri et al., 2016)(Brizzi et al., 2018)(Bian et al., 2018)
Medical Robotics
(Zinchenko and Song, 2021)(Trejo and Hu, 2018)(Black et al., 2023)
Other
(Mourtzis et al., 2019)(Mourtzis et al., 2017)(Yew et al., 2017)(Xie et al., 2022)(Mu et al., 2021)(Sun et al., 2021)(Zaman et al., 2023)(Liu et al., 2023)
Robotic arm+mobile robot
(Chen et al., 2020)(Abdulsalam and Aouf, 2023)(Du et al., 2022a)

### v3

@@ Abstract.
The rising interest of generalist robots seek to create robots with versatility to handle multiple tasks in a variety of environments, and human will interact with such robots through immersive interfaces. In the context of human-robot interaction (HRI), this survey provides an exhaustive review of the applications of extended reality (XR) technologies in the field of remote HRI. We developed a systematic search strategy based on the PRISMA methodology. From the initial 2,561 articles selected, 100 research papers that met our inclusion criteria were included. We categorized and summarized the domain in detail, delving into XR technologies, including augmented reality (AR), virtual reality (VR), and mixed reality (MR), and their applications in facilitating intuitive and effective remote control and interaction with robotic systems. The survey highlights existing articles on the application of XR technologies, user experience enhancement, and various interaction designs for XR in remote HRI, providing insights into current trends and future directions. We also identified potential gaps and opportunities for future research to improve remote HRI systems through XR technology to guide and inform future XR and robotics research.
@@ 2.2. Inclusion and Exclusion Criteria
Criterion Description
$I_{1}$ The research proposes the design, development or system for remote control of robots based on augmented, virtual or mixed reality.
$I_{2}$ The research used augmented, virtual or mixed reality as a method of remote control of the robot.
$E_{1}$ The system proposed in the study does not support remote control.
$E_{2}$ The study does not involve real robots.
$E_{3}$ The target of research control is not a robot.
$E_{4}$ The study did not use augmented, virtual or mixed reality technologies.
$E_{5}$ This is not a technical article.
@@ 2.3. Data Extraction and Analysis
ID Data Extraction Type
DE1 Study ID Open text
DE2 Title Open text
DE3 Keywords Open text
DE4 Used XR technologies VR HMD, AR HMD, MR, Mobile AR, CAVE, Other
DE5 Interaction modalities Gesture, Controller, Joystick, Gaze, Head, Haptic devices, Motion capture, Walk, 2D screen, Voice, Glove, Other
DE6 Virtual interfaces Direct, Digital twin, Virtual control room, Digital twin+3D reconstruction, Direct+3D reconstruction, 3D reconstruction, Context-aware AR, Multiple, Other
DE7 User perspective Coupling with robots, Decoupling from robots, Dynamic perspective, Bird’s-eye view, Other
DE8 Type of robots Mobile robot, Drones/UAV, Humanoid Robot, Robotic arm, Mobile robot+robotic arm, Double-armed robot, Medical Robotics, Other
DE9 Specific tasks involved Navigation, Grabbing/Picking/Placement, Surgery, Game/Entertainment, Industrial/Manufacturing, Search, Environment scan, No, Multiple, Other
DE10 How was it measured or evaluated? Time/accuracy of the task, Interviews, Questionnaire, AR/VR performance, Comparison, N/A, Other
DE11 Was there a discussion about delays? No, Yes(Times), Other
DE12 Support multiplayer collaboration? No, Multi-user - one robot, One user - multi-robot, Multi - Multi, Other
DE13 Where are the enhancements located? User, Robot (real), Robot (virtual), Object (virtual), Real environment(RE), Virtual environment(VE)
DE14 Enhancement types Haptic, Voice, Graphic, Text, 3D Object, Highlight, Ray, Avatar
@@ Appendix A Data Extraction List for Included Publication
VR HMD (Chen et al., 2022)(Xu et al., 2022b)(Ponomareva et al., 2021)(Wei et al., 2021)(Vu et al., 2022)(Lipton et al., 2017)(Su et al., 2021)(Walker et al., 2021)(Wibowo et al., 2021)(Zein et al., 2021)(Vagvolgyi et al., 2018)(Kulikovskiy et al., 2021)(Zhou et al., 2020)(Chen et al., 2017)(Park et al., 2018)(Stotko et al., 2019)(Peppoloni et al., 2015b)(Wang et al., 2019a)(Chen et al., 2020)(Galambos et al., 2015)(Kuo et al., 2021)(Yun and Jun, 2022)(Xie et al., 2022)(Schwarz et al., 2021)(Cichon and Roßmann, 2018)(Vempati et al., 2019)(Fani et al., 2018b)(Zinchenko and Song, 2021)(Codd-Downey et al., 2014)(Bai et al., 2018)(Peppoloni et al., 2015a)
@@
AR HMD (Mourtzis et al., 2019)(Hedayati et al., 2018)(Mourtzis et al., 2017)(Xie et al., 2022)(Hernández et al., 2020)(Gai et al., 2020)(Du et al., 2022b)(Walker et al., 2019)(Xue Er Shamaine et al., 2020)(Du et al., 2019)(Choi et al., 2022)(Gong et al., 2017)(Xue et al., 2020)(Aschenbrenner et al., 2018)(Szczurek et al., 2023)(Steinke et al., 2023)(Meng et al., 2023)(Rivera-Pinto et al., 2023)(Black et al., 2023)(Du et al., 2022a)
MR (Cruz Ulloa et al., 2022)(Ai et al., 2016)(Hönig et al., 2015)(Ihara et al., 2023)(Zaman et al., 2023)
CAVE (Galambos et al., 2015)(Gammieri et al., 2017)(Betancourt et al., 2022)
Mobie AR (Chacko et al., 2020)
Other (Stedman et al., 2022)(Rastogi and Kumar Srivastava, 2019)(Yew et al., 2017)(Mu et al., 2021)
@@
Controller (Chen et al., 2022)(Wei et al., 2021)(Lipton et al., 2017)(Walker et al., 2021)(Kulikovskiy et al., 2021)(Zhou et al., 2020)(Wang et al., 2019a)(Chen et al., 2020)(Gammieri et al., 2017)(Kuo et al., 2021)(Vempati et al., 2019)(Codd-Downey et al., 2014)(Takacs et al., 2015)(Omarali et al., 2020)(Xu et al., 2022a)(Kalinov et al., 2021)(Hernández et al., 2020)(Le et al., 2020)(Wang et al., 2017)(Nagy et al., 2019)(Nandhini et al., 2023)(Abdulsalam and Aouf, 2023)(Kanazawa et al., 2023)(Fuchino et al., 2023)(Nenna et al., 2023)(Sanfilippo et al., 2022)(Yun and Jun, 2022)(Zaman et al., 2023)(Fan et al., 2023)(Liu et al., 2023)
Gesture (Cruz Ulloa et al., 2022)(Mourtzis et al., 2017)(Xie et al., 2022)(Peppoloni et al., 2015b)(Jang et al., 2021a)(Jang et al., 2021b)(Zhao et al., 2017)(Xu et al., 2018)(Cousins et al., 2017)(Xue Er Shamaine et al., 2020)(Bian et al., 2018)(Choi et al., 2022)(Xue et al., 2020)(Ihara et al., 2023)(Szczurek et al., 2023)(Meng et al., 2023)(Black et al., 2023)(Du et al., 2022b)(Sakashita et al., 2023)(Zaman et al., 2023)(Rivera-Pinto et al., 2023)
Joystick (Vu et al., 2022)(Wibowo et al., 2021)(Zein et al., 2021)(Hedayati et al., 2018)(Stotko et al., 2019)(Ai et al., 2016)(Betancourt et al., 2022)(Walker et al., 2019)(Xia et al., 2023)
Haptic devices (Ponomareva et al., 2021)(Chen et al., 2017)(Su et al., 2021)(Rastogi and Kumar Srivastava, 2019)(Schwarz et al., 2021)(Trejo and Hu, 2018)(Kim et al., 2023)(Yun and Jun, 2022)(Fan et al., 2023)
Motion capture (Xu et al., 2022b)(Peppoloni et al., 2015b)(Fani et al., 2018b)(Bai et al., 2018)(Brizzi et al., 2018)
Walk (Phan et al., 2018)(Gai et al., 2020)(Grzeskowiak et al., 2020)(Du et al., 2019)(Sakashita et al., 2023)
2D screen (Mu et al., 2021)(Chacko et al., 2020)(Gong et al., 2017)
Voice (Hormaza et al., 2019)(Aschenbrenner et al., 2018)(Rivera-Pinto et al., 2023)
Glove (Park et al., 2018)(Sun et al., 2021)(Trinitatova and Tsetserukou, 2023)
Head (Theofilis et al., 2016)(Sakashita et al., 2023)
Gaze (Moniri et al., 2016)
Other (Stedman et al., 2022)(Mourtzis et al., 2019)(Vagvolgyi et al., 2018)(Galambos et al., 2015)(Cichon and Roßmann, 2018)(Zinchenko and Song, 2021)(Wang et al., 2019c)(Hönig et al., 2015)(Wang et al., 2021b)(Du et al., 2019)(McHenry et al., 2021)(Liu et al., 2023)
@@
Digital twin (Cruz Ulloa et al., 2022)(Su et al., 2021)(Vagvolgyi et al., 2018)(Wang et al., 2019a)(Yew et al., 2017)(Gammieri et al., 2017)(Yun and Jun, 2022)(Xie et al., 2022)(Zinchenko and Song, 2021)(Peppoloni et al., 2015b)(Takacs et al., 2015)(Jang et al., 2021b)(Trejo and Hu, 2018)(Betancourt et al., 2022)(Le et al., 2020)(Wang et al., 2021b)(Du et al., 2022b)(Walker et al., 2019)(Brizzi et al., 2018)(Xue Er Shamaine et al., 2020)(Grzeskowiak et al., 2020)(Wang et al., 2019b)(Nagy et al., 2019)(Choi et al., 2022)(McHenry et al., 2021)(Xue et al., 2020)(Nandhini et al., 2023)(Fan et al., 2023)(Abdulsalam and Aouf, 2023)(Trinitatova and Tsetserukou, 2023)(Meng et al., 2023)
@@
Direct (Xu et al., 2022b)(Wibowo et al., 2021)(Hedayati et al., 2018)(Hedayati et al., 2018)(Chen et al., 2017)(Park et al., 2018)(Stotko et al., 2019)(Peppoloni et al., 2015b)(Chen et al., 2020)(Ai et al., 2016)(Schwarz et al., 2021)(Fani et al., 2018b)(Bai et al., 2018)(Wang et al., 2019c)(Mu et al., 2021)(Theofilis et al., 2016)(Zhao et al., 2017)(Gai et al., 2020)(Cousins et al., 2017)(Chacko et al., 2020)(Brizzi et al., 2018)(Bian et al., 2018)(Zhang, 2018)(Gong et al., 2017)(Ihara et al., 2023)(Zaman et al., 2023)(Xia et al., 2023)(Black et al., 2023)(Sanfilippo et al., 2022)
Digital twin+3D reconstruction (Ponomareva et al., 2021)(Stedman et al., 2022)(Rastogi and Kumar Srivastava, 2019)(Galambos et al., 2015)(Kuo et al., 2021)(Cichon and Roßmann, 2018)(Codd-Downey et al., 2014)(Phan et al., 2018)(Omarali et al., 2020)(Jang et al., 2021a)(Hönig et al., 2015)(Wang et al., 2017)(Szczurek et al., 2023)(Du et al., 2022b)(Aschenbrenner et al., 2018)
Multiple (Vu et al., 2022)(Zhou et al., 2020)(Xu et al., 2022a)(Du et al., 2019)(Kanazawa et al., 2023)(Liu et al., 2023)
Direct+3D reconstruction (Chen et al., 2022)(Zein et al., 2021)(Vempati et al., 2019)(Sakashita et al., 2023)(Kim et al., 2023)
3D reconstruction (Wei et al., 2021)(Walker et al., 2021)(Moniri et al., 2016)
Virtual control room (Lipton et al., 2017)(Kulikovskiy et al., 2021)(Steinke et al., 2023)
Other (Mourtzis et al., 2019)(Mourtzis et al., 2017)(Hernández et al., 2020)(Sun et al., 2021)(Xu et al., 2018)(Hormaza et al., 2019)(Fuchino et al., 2023)
@@
Decoupling with robots (Ponomareva et al., 2021)(Vu et al., 2022)(Stedman et al., 2022)(Cruz Ulloa et al., 2022)(Mourtzis et al., 2019)(Su et al., 2021)(Hedayati et al., 2018)(Vagvolgyi et al., 2018)(Wang et al., 2019a)(Rastogi and Kumar Srivastava, 2019)(Yew et al., 2017)(Galambos et al., 2015)(Gammieri et al., 2017)(Kuo et al., 2021)(Yun and Jun, 2022)(Xie et al., 2022)(Cichon and Roßmann, 2018)(Zinchenko and Song, 2021)(Codd-Downey et al., 2014)(Peppoloni et al., 2015a)(Omarali et al., 2020)(Jang et al., 2021b)(Trejo and Hu, 2018)(Moniri et al., 2016)(Xu et al., 2018)(Le et al., 2020)(Hönig et al., 2015)(Wang et al., 2021b)(Du et al., 2022b)(Hormaza et al., 2019)(Walker et al., 2019)
@@
Coupling with robots (Xu et al., 2022b)(Lipton et al., 2017)(Wibowo et al., 2021)(Zein et al., 2021)(Kulikovskiy et al., 2021)(Chen et al., 2017)(Park et al., 2018)(Stotko et al., 2019)(Peppoloni et al., 2015b)(Mourtzis et al., 2017)(Chen et al., 2020)(Ai et al., 2016)(Schwarz et al., 2021)(Vempati et al., 2019)(Fani et al., 2018b)(Bai et al., 2018)(Wang et al., 2019c)(Mu et al., 2021)(Kalinov et al., 2021)(Theofilis et al., 2016)(Zhao et al., 2017)(Gai et al., 2020)(Sun et al., 2021)(Cousins et al., 2017)(Chacko et al., 2020)(Brizzi et al., 2018)(Bian et al., 2018)(Zhang, 2018)(Gong et al., 2017)(Sakashita et al., 2023)(Kim et al., 2023)(Steinke et al., 2023)
@@
Dynamic perspective (Zhou et al., 2020)(Xu et al., 2022a)(Hernández et al., 2020)(Betancourt et al., 2022)(Liu et al., 2023)
God perspective (Walker et al., 2021)(Jang et al., 2021a)
Other (Chen et al., 2022)(Vu et al., 2022)(Takacs et al., 2015)(Phan et al., 2018)(Grzeskowiak et al., 2020)(Aschenbrenner et al., 2018)(Fuchino et al., 2023)(Black et al., 2023)
@@
Robotic arm (Xu et al., 2022b)(Ponomareva et al., 2021)(Wei et al., 2021)(Vu et al., 2022)(Su et al., 2021)(Vagvolgyi et al., 2018)(Chen et al., 2017)(Wang et al., 2019a)(Rastogi and Kumar Srivastava, 2019)(Galambos et al., 2015)(Gammieri et al., 2017)(Kuo et al., 2021)(Yun and Jun, 2022)(Fani et al., 2018b)(Bai et al., 2018)(Peppoloni et al., 2015a)(Omarali et al., 2020)(Xu et al., 2022a)(Jang et al., 2021b)(Xu et al., 2018)(Le et al., 2020)(Cousins et al., 2017)(Wang et al., 2021b)(Hormaza et al., 2019)(Xue Er Shamaine et al., 2020)(Wang et al., 2019b)(Du et al., 2019)(Nagy et al., 2019)(Choi et al., 2022)(McHenry et al., 2021)
@@
Mobile robot (Stedman et al., 2022)(Cruz Ulloa et al., 2022)(Walker et al., 2021)(Wibowo et al., 2021)(Stotko et al., 2019)(Cichon and Roßmann, 2018)(Codd-Downey et al., 2014)(Jang et al., 2021a)(Hernández et al., 2020)(Zhao et al., 2017)(Gai et al., 2020)(Du et al., 2022b)(Chacko et al., 2020)(Wang et al., 2017)(Zhang, 2018)(Ihara et al., 2023)(Sakashita et al., 2023)(Szczurek et al., 2023)(Kanazawa et al., 2023)(Xia et al., 2023)
Drones/UAV (Zein et al., 2021)(Hedayati et al., 2018)(Ai et al., 2016)(Vempati et al., 2019)(Takacs et al., 2015)(Wang et al., 2019c)(Kalinov et al., 2021)(Betancourt et al., 2022)(Hönig et al., 2015)(Walker et al., 2019)
Humanoid robot (Chen et al., 2022)(Kulikovskiy et al., 2021)(Schwarz et al., 2021)(Takacs et al., 2015)(Theofilis et al., 2016)(Grzeskowiak et al., 2020)(Gong et al., 2017)(Sanfilippo et al., 2022)
Double-armed robot (Lipton et al., 2017)(Zhou et al., 2020)(Park et al., 2018)(Peppoloni et al., 2015b)(Moniri et al., 2016)(Brizzi et al., 2018)(Bian et al., 2018)
Medical Robotics (Zinchenko and Song, 2021)(Trejo and Hu, 2018)(Black et al., 2023)
Other (Mourtzis et al., 2019)(Mourtzis et al., 2017)(Yew et al., 2017)(Xie et al., 2022)(Mu et al., 2021)(Sun et al., 2021)(Zaman et al., 2023)(Liu et al., 2023)
Robotic arm+mobile robot (Chen et al., 2020)(Abdulsalam and Aouf, 2023)(Du et al., 2022a)